THE MILLIMETER/SUBMILLIMETER SPECTRUM OF CUSH(X̃^{1}A′): INVESTIGATION OF THE TRAN-SITION METAL SULFUR BOND

\underline{A. J. JANCZYK} and L. M. ZIURYS, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ, 85721.

The pure rotational spectra of ^{63}CuSH and ^{65}CuSH and its deuterated isotopomers were recorded using millimeter/submillimeter-wave direct absorption techniques. The molecule was produced by reacting copper vapor with H_{2}S in the presence of d.c. discharge. Transitions ranging from J = 24 → 25 to J = 50 → 51 spanning the region of 262 - 538 GHz were measured for ^{63}CuSH and ^{65}CuSH, as well as transitions J = 32 → 33 to J = 39 → 40 for the deuterated species. The spectra showed extensive K_{a} ladder structure for all species, which is consistent with the molecule being an asymmetric top. Rotational constants, as well as an r_{0} structure, have been determined for the molecule. This study shows that the molecule is bent with a Cu-S-H angle of Θ = 93.6^{o}, and therefore is different from CuOH, which has a Cu-O-H angle of Θ =110.1^{o}. This striking variation in geometry likely results from a lack of hybridization in the CuSH molecule.